Can I have the answer please

answer in the form of a question, to the statement below

1-"The length of an arrow is used to represent this, in a vector"

2.An airplane making an approach for landing decelerates at a rate of 2m/s2. How long will it take the airplane to reduce its speed by 264km/h
Answer #1

ANSWER: Length of an arrow is used to represent magnitude in a vector

------------------------------------

deceleration =2m/s2

Given reduction in speed as 264km/hr

Convert it into m/s

264km/hr-73.3m/s

so initial speed - final speed =73.3 m/s

73 3 u - u-m/s

Use Formula v=u+at

v-u=at

u-v=-at

73 2

73.3 2 * t

ANSWER: t 36.65s

===================